### v1

@@ 5.2 Combining Round Robin and Bag-Filling
Note that the input $\mathcal{I}$ of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ is $R^{\alpha}$-irreducible. For all the agents $i$, we define $\mu_{i}=\mu^{n}_{i}(M)$ to be the MMS value of agent $i$ in the beginning of the Algorithm 4. We do a round robin fashion process as following: starting from $n$ empty bags $B_{1},\ldots,B_{n}$, for $j$ in the cyclic sequence $1,2,\ldots,n$, we add the most valuable available good (not yet allocated to any bag) to $B_{j}$. If for some remaining agent $i$, $v_{i}(B_{j})\geq\alpha\cdot\mu_{i}$, then we allocate $B_{j}$ to $i$ and remove $i$ from the set of agents and $j$ from the cyclic sequence. We call this phase of the algorithm, the round robin phase. Note that for all the remaining bags $B_{a}$ and $B_{b}$, if $a<b$, then $v_{i}(B_{a})\geq v_{i}(B_{b})$ for all agents $i$ and hence, first index $1$ will be removed, then index $2$ and so on (see Lemma 14). Let $\{B_{a},\ldots,B_{n}\}$ be the set of the remaining bags at the end of the round robin phase. In the next phase of the algorithm, called duplication phase, we duplicate goods $11,2,\ldots,n-a+$ 1, add them to $B_{a},\ldots,B_{n}$ respectively and allocate these bags to the remaining agents arbitrarily. Unlike the round robin phase, it is not clear whether the agents who are allocated during the duplication phase value their bundle at least $\alpha$ fraction of their MMS. However, in Sections 5.3 and 5.4, we show that by appropriately reducing the instance in advance, we can ensure both the desired approximation guarantees and the bound on the number of distinct copies.
@@
For the rest of this (sub-)section, let $\alpha\geq 0$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible instance, and $\mu_{i}$ be $\mu_{i}(\mathcal{I})$.
@@
For all the remaining agents $i$ after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, and all the bags $B_{j}$ with $|B_{j}|\geq 3$ that were allocated to other agents during the round robin phase, we have $v_{i}(B_{j})\leq(4\alpha/3)\cdot\mu_{i}$.
@@
Let $g$ be the last good added to $B_{j}$. We have $g\geq 2n+1$ and hence $v_{i}(g)\leq v_{i}(2n+1)\leq\alpha\mu_{i}/3$. The last inequality holds since the instance is $R^{\alpha}_{2}$-irreducible. We have $v_{i}(B_{j}\setminus\{g\})<\alpha\cdot\mu_{i}$. Otherwise, $B_{j}\setminus\{g\}$ would have been allocated to $i$, or some other agent who values it at least $\alpha$ fraction of their MMS value, previously. Therefore we have,
$v_{i}(B_{j})$ $=v_{i}(B_{j}\setminus\{g\})+v_{i}(g)\leq\alpha\cdot\mu_{i}+\alpha/3\cdot\mu_{i}=4\alpha/3\cdot\mu_{i}.\qed$
@@
We prove that first $B_{1}$ gets allocated, then $B_{2}$ and so on. Towards contradiction, assume for some $j>1$ bag $B_{j}$ gets allocated to agent $i$ while bag $B_{j-1}$ is not assigned. At the time that $B_{j}$ gets assigned, we have $|B_{j-1}|=|B_{j}|$ since both $B_{j-1}$ and $B_{j}$ were present in all the rounds so far. Let $B_{j-1}=\{g_{1},\ldots,g_{k}\}$ and $B_{j}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k}\}$. Followed from round robin, for all $\ell\in[k]$, we have $g_{\ell}=g^{\prime}_{\ell}-1$ which means $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell})$. Hence $v_{i}(B_{j-1})\geq v_{i}(B_{j})\geq\mu_{i}$ which is a contradiction that $B_{j-1}$ was not assigned to anyone before moving to $B_{j}$. ∎
@@ 5.3 $6/7$-MMS with Copies
In this section, we present an algorithm which given an ordered instance, outputs an $\alpha$-MMS allocation with at most $\lfloor n/2\rfloor$ distinct copies for any $\alpha\leq 6/7$.
First we reduce the instance in Algorithm 5 using $R^{\alpha}_{0},R^{\alpha}_{1},\ldots,$ and $S^{\alpha}$. It is not difficult to see that if $k$ agents are removed (i.e. are assigned a bundle) by the end of the algorithm, then in total at most $\lfloor k/2\rfloor$ duplicated goods are introduced and each good is duplicated at most once. Also, since we remove these $k$ agents and all the goods in these $k$ bags, in the future no more copies of them can be used. Since all these rules are $\alpha$-valid reductions, it suffices to give an $\alpha$-MMS allocation with at most $\lfloor(n^{*}-k)/2\rfloor$ distinct copies for the instance obtained by the remaining agents and goods where $n^{*}$ is the original number of agents. Let $n=n^{*}-k$.
@@
For the rest of this (sub)-section, let $\alpha\leq 6/7$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible and $S^{\alpha}$-irreducible instance, and $\mu_{i}$ be $\mu_{i}(\mathcal{I})$. Let $n$ be the number of agents right before the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, =$\mathcal{B}=\{B_{1},\ldots,B_{a-1}\}$ be the set of bags allocated during this phase, and $n^{\prime}=n-a+1$ be the number of remaining agents after this phase.
Observation 1.
@@
Let $i$ be a remaining agent after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ (if any). Then |$|\{B\in\mathcal{B}\mid v_{i}(B)>\mu_{i}\}|\geq n^{\prime}$.
@@
Let +=$\mathcal{B}^{+}=\{B\in\mathcal{B}\mid v_{i}(B)>\mu_{i}\}$. By definition, for all $B\in\mathcal{B}\setminus\mathcal{B}^{+}$, $v_{i}(B)\leq\mu_{i}$. By Observation 1, for all $B\in\mathcal{B}$ , $|B|\geq 3$. By Lemma 13, for all $B\in\mathcal{B}^{+}$, $v_{i}(B)\leq 4\alpha/3\cdot\mu_{i}$. Also, let $\bar{\mathcal{B}}$ be the set of $n-|\mathcal{B}|=n^{\prime}$ remaining bags after the round robin phase. For all $B\in\bar{\mathcal{B}}$, $v_{i}(B)<\alpha\cdot\mu_{i}$. Otherwise, $B$ should have been allocated to $i$ (or some other agent who values it at least $\alpha$ fraction of their MMS value). Therefore, we have
$n\cdot\mu_{i}$ $\leq v_{i}(M)$
@@
$\leq|\mathcal{B}^{+}|4\alpha/3\cdot\mu_{i}+(|\mathcal{B}|-|\mathcal{B}^{+}|)\mu_{i}+(n-|\mathcal{B}|)\alpha\cdot\mu_{i}$
$\leq(|\mathcal{B}|-|\mathcal{B}^{+}|)\mu_{i}+6/7(|\mathcal{B}^{+}|/3+n-|\mathcal{B}|+|\mathcal{B^{+}}|)\cdot\mu_{i}$ ($\alpha\leq 6/7$)
Dividing both sides of the inequality by $\mu_{i}$ and removing $(|\mathcal{B}|-|\mathcal{B}^{+}|)$, we obtain,
$n-|\mathcal{B}|+|\mathcal{B}^{+}|$ $\leq 6/7(|\mathcal{B}^{+}|/3+n-|\mathcal{B}|+|\mathcal{B}^{+}|)=2|\mathcal{B}^{+}|/7+6/7(n-|\mathcal{B}|+|\mathcal{B}^{+}|).$
Therefore, $|\mathcal{B}^{+}|\geq(n-|\mathcal{B}|+|\mathcal{B}^{+}|)/2$ which means $|\mathcal{B}^{+}|\geq n-|\mathcal{B}|=n^{\prime}$. ∎
@@
For all the remaining agents $i$ after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, and all the remaining bags $B_{j}$, we have $v_{i}(B_{j}\cup\{n^{\prime}\})\geq\mu_{i}$.
@@
By Lemma 14, the bags $B_{1},\ldots,B_{n-n^{\prime}}$ are allocated during the round robin phase. By Lemma 16, there exists at least $n^{\prime}$ many bags $B_{i_{1}},\ldots,B_{i_{n^{\prime}}}$ for $i_{1}<i_{2}<\ldots<i_{n^{\prime}}\leq n-n^{\prime}$ that were allocated to other agents during the round robin phase and $v_{i}(B_{i_{\ell}})>\mu_{i}$ for all $\ell\in[n^{\prime}]$. Note that $i_{n^{\prime}}\geq n^{\prime}$. Since $B_{j}$ have been assigned a good in all but maybe the last round, we have $|B_{j}|\geq|B_{i_{n^{\prime}}}|-1$. Let $B_{j}=\{g_{1},\ldots,g_{k}\}$ and $B_{i_{n^{\prime}}}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k^{\prime}}\}$ for $j=g_{1}<\ldots<g_{k}$ and $i_{n^{\prime}}=g^{\prime}_{1}<\ldots<g^{\prime}_{k^{\prime}}$. For all $\ell\in[k^{\prime}-1]$, since $g_{\ell}$ is allocated before $g^{\prime}_{\ell+1}$, $g_{\ell}<g^{\prime}_{\ell+1}$ and hence $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell+1})$. We have
@@
$\geq\sum_{\ell\in[k^{\prime}-1]}v_{i}(g^{\prime}_{\ell+1})+v_{i}(i_{n^{\prime}})=v_{i}(B_{i_{n^{\prime}}})>\mu_{i}.\qed$
@@
returns an $\alpha$-MMS allocation with at most $\lfloor n^{*}/2\rfloor$ many distinct copies.
@@
Let $n$ be the number of agents after applying $\mathtt{S-reduce(\mathcal{I},\alpha)}$. Since $S^{\alpha}$ and $R^{\alpha}_{k}$ are valid reductions for all $k<|M|/n^{*}$, all agents who receive a bag during the $\mathtt{S-reduce(\mathcal{I},\alpha)}$ receive at least $\alpha$ fraction of their MMS value and the MMS value of the remaining agents do not decrease after this process. All the agents who receive a bag during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}$ get at least $\alpha$ fraction of their value. Let agent $i$ be an agent who receives bag $B_{j}$ after the round robin phase. Let $g$ be the good that was duplicated and added to $B_{j}$ right before allocating it to agent $i$. $g\leq n^{\prime}$ and thus $v_{i}(g)\geq v_{i}(n^{\prime})$. By Lemma 17, $v_{i}(B_{j}\cup\{g\})\geq\alpha\mu_{i}$. Therefore, the final allocation is $\alpha$-MMS.
Now we prove at most $\lfloor n^{*}/2\rfloor$ distinct copies are used. Let $k=n^{*}-n$ be the number of agents who receive a bag in $\mathtt{S-reduce(\mathcal{I},\alpha)}$. Only when applying $S^{\alpha}$, new copies are introduced and one copy per two agents that are removed. Furthermore, the goods that are allocated during $S^{\alpha}$ along with their copies are removed from the instance. Therefore, they cannot be copied more than once. So during $\mathtt{S-reduce(\mathcal{I},\alpha)}$, at most $\lfloor k/2\rfloor$ distinct copies are used. In the duplication phase of $\mathsf{BagFill\mbox{-}RoundRobin}$, we duplicate items $1,2,\ldots,n^{\prime}$ once. By Corollary 6, $n^{\prime}\leq\lfloor n/2\rfloor$. Therefore, in total at most $\lfloor n^{*}/2\rfloor$ distinct copies are used. ∎

### v2

@@ 5.3 Combining Round Robin and Bag-Filling
Note that the input $\mathcal{I}$ of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ is $R^{\alpha}$-irreducible. For all the agents $i$, we define $\mu_{i}=\mu^{n}_{i}(M)$ to be the MMS value of agent $i$ in the beginning of the Algorithm 4. We do a round robin fashion process as following: starting from $n$ empty bags $B_{1},\ldots,B_{n}$, for $j$ in the cyclic sequence $1,2,\ldots,n$, we add the most valuable available good (not yet allocated to any bag) to $B_{j}$. If for some remaining agent $i$, $v_{i}(B_{j})\geq\alpha\cdot\mu_{i}$, then we allocate $B_{j}$ to $i$ and remove $i$ from the set of agents and $j$ from the cyclic sequence. We call this phase of the algorithm, the round robin phase. Note that for all the remaining bags $B_{a}$ and $B_{b}$, if $a<b$, then $v_{i}(B_{a})\geq v_{i}(B_{b})$ for all agents $i$ and hence, first index $1$ will be removed, then index $2$ and so on (see Lemma 15). Let $\{B_{a},\ldots,B_{n}\}$ be the set of the remaining bags at the end of the round robin phase. In the next phase of the algorithm, called duplication phase, we duplicate goods $11,2,\ldots,n-a+$ 1, add them to $B_{a},\ldots,B_{n}$ respectively and allocate these bags to the remaining agents arbitrarily. Unlike the round robin phase, it is not clear whether the agents who are allocated during the duplication phase value their bundle at least $\alpha$ fraction of their MMS. However, in Sections 5.4 and 5.5, we show that by appropriately reducing the instance in advance, we can ensure both the desired approximation guarantees and the bound on the number of distinct copies.
@@
For the rest of this (sub-)section, let $\alpha\geq 0$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible instance, and $\mu_{i}(\mathcal{I})=1$ without loss of generality.
@@
For all the remaining agents $i$ after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, and all the bags $B_{j}$ with $|B_{j}|\geq 3$ that were allocated to other agents during the round robin phase, we have $v_{i}(B_{j})\leq 4\alpha/3$.
@@
Let $g$ be the last good added to $B_{j}$. We have $g\geq 2n+1$ and hence $v_{i}(g)\leq v_{i}(2n+1)\leq\alpha/3$. The last inequality holds since the instance is $R^{\alpha}_{2}$-irreducible (by Observation 1). We have $v_{i}(B_{j}\setminus\{g\})<\alpha$. Otherwise, $B_{j}\setminus\{g\}$ would have been allocated to $i$, or some other agent who values it at least $\alpha$, previously. Therefore we have,
$v_{i}(B_{j})$ $=v_{i}(B_{j}\setminus\{g\})+v_{i}(g)\leq\alpha+\alpha/3=4\alpha/3.\qed$
@@
We prove that first $B_{1}$ gets allocated, then $B_{2}$ and so on. Towards contradiction, assume for some $j>1$ bag $B_{j}$ gets allocated to agent $i$ while bag $B_{j-1}$ is not assigned. At the time that $B_{j}$ gets assigned, we have $|B_{j-1}|=|B_{j}|$ since both $B_{j-1}$ and $B_{j}$ were present in all the rounds so far. Let $B_{j-1}=\{g_{1},\ldots,g_{k}\}$ and $B_{j}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k}\}$. Followed from round robin, for all $\ell\in[k]$, we have $g_{\ell}=g^{\prime}_{\ell}-1$ which means $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell})$. Hence $v_{i}(B_{j-1})\geq v_{i}(B_{j})\geq 1$ which is a contradiction that $B_{j-1}$ was not assigned to anyone before moving to $B_{j}$. ∎
@@ 5.4 $6/7$-MMS with Copies
In this section, we present an algorithm which given an ordered instance, outputs a simple $\alpha$-MMS allocation with at most $\lfloor n/2\rfloor$ distinct copies for any $\alpha\leq 6/7$.
First we reduce the instance in Algorithm 5 using $R^{\alpha}_{0},R^{\alpha}_{1},\ldots,$ and $S^{\alpha}$. It is not difficult to see that if $k$ agents are removed (i.e. are assigned a bundle) by the end of the algorithm, then in total at most $\lfloor k/2\rfloor$ duplicated goods are introduced, each good is duplicated at most once, and all the instances of these goods (original and copy) are allocated to different agents. Also, since we remove these $k$ agents and all the goods in these $k$ bags, in the future no more copies of them can be used. Since all these rules are $\alpha$-valid reductions, it suffices to give a simple $\alpha$-MMS allocation with at most $\lfloor(n^{*}-k)/2\rfloor$ distinct copies for the instance obtained by the remaining agents and goods where $n^{*}$ is the original number of agents. Let $n=n^{*}-k$.
@@
For the rest of this (sub)-section, let $\alpha\leq 6/7$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible and $S^{\alpha}$-irreducible instance, and $\mu_{i}(\mathcal{I})=1$. Let $n$ be the number of agents right before the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, =$\mathcal{B}=\{B_{1},\ldots,B_{a-1}\}$ be the set of bags allocated during this phase, and $n^{\prime}=n-a+1$ be the number of remaining agents after this phase.
Observation 2.
@@
Proof.
Since $\mathcal{I}$ is $S^{\alpha}$-irreducible, $v_{i}(\{j,n+j\})\leq v_{i}(\{1,n+1\})<\alpha$ for all $i\in N$. Thus, if $B_{j}$ was allocated to some agent, $|B_{j}|>2$. ∎
@@
Let $i$ be a remaining agent after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ (if any). Then |$|\{B\in\mathcal{B}\mid v_{i}(B)>1\}|>n^{\prime}$.
@@
Let +=$\mathcal{B}^{+}=\{B\in\mathcal{B}\mid v_{i}(B)>1\}$. By definition, for all $B\in\mathcal{B}\setminus\mathcal{B}^{+}$, $v_{i}(B)\leq 1$. By Observation 2, for all $B\in\mathcal{B}$ , $|B|\geq 3$. By Lemma 14, for all $B\in\mathcal{B}^{+}$, $v_{i}(B)\leq 4\alpha/3$. Also, let $\bar{\mathcal{B}}$ be the set of $n-|\mathcal{B}|=n^{\prime}$ remaining bags after the round robin phase. For all $B\in\bar{\mathcal{B}}$, $v_{i}(B)<\alpha$. Otherwise, $B$ should have been allocated to $i$ (or some other agent who values it at least $\alpha$). Therefore, we have
$n$ $\leq v_{i}(M)$
@@
$<|\mathcal{B}^{+}|4\alpha/3+(|\mathcal{B}|-|\mathcal{B}^{+}|)+(n-|\mathcal{B}|)\alpha$
$\leq(|\mathcal{B}|-|\mathcal{B}^{+}|)+6/7(|\mathcal{B}^{+}|/3+n-|\mathcal{B}|+|\mathcal{B^{+}}|)$ ($\alpha\leq 6/7$)
Removing $(|\mathcal{B}|-|\mathcal{B}^{+}|)$ from both sides of the inequality, we obtain,
$n-|\mathcal{B}|+|\mathcal{B}^{+}|$ $<6/7(|\mathcal{B}^{+}|/3+n-|\mathcal{B}|+|\mathcal{B}^{+}|)=2|\mathcal{B}^{+}|/7+6/7(n-|\mathcal{B}|+|\mathcal{B}^{+}|).$
Therefore, $|\mathcal{B}^{+}|>(n-|\mathcal{B}|+|\mathcal{B}^{+}|)/2$ which means $|\mathcal{B}^{+}|>n-|\mathcal{B}|=n^{\prime}$. ∎
@@
For all the remaining agents $i$ after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, and all the remaining bags $B_{j}$, we have $v_{i}(B_{j}\cup\{n^{\prime}\})\geq 1$.
@@
By Lemma 15, the bags $B_{1},\ldots,B_{n-n^{\prime}}$ are allocated during the round robin phase. By Lemma 17, there exists at least $n^{\prime}$ many bags $B_{i_{1}},\ldots,B_{i_{n^{\prime}}}$ for $i_{1}<i_{2}<\ldots<i_{n^{\prime}}\leq n-n^{\prime}$ that were allocated to other agents during the round robin phase and $v_{i}(B_{i_{\ell}})>1$ for all $\ell\in[n^{\prime}]$. Note that $i_{n^{\prime}}\geq n^{\prime}$. Since $B_{j}$ have been assigned a good in all but maybe the last round, we have $|B_{j}|\geq|B_{i_{n^{\prime}}}|-1$. Let $B_{j}=\{g_{1},\ldots,g_{k}\}$ and $B_{i_{n^{\prime}}}=\{g^{\prime}_{1},\ldots,g^{\prime}_{k^{\prime}}\}$ for $j=g_{1}<\ldots<g_{k}$ and $i_{n^{\prime}}=g^{\prime}_{1}<\ldots<g^{\prime}_{k^{\prime}}$. For all $\ell\in[k^{\prime}-1]$, since $g_{\ell}$ is allocated before $g^{\prime}_{\ell+1}$, $g_{\ell}<g^{\prime}_{\ell+1}$ and hence $v_{i}(g_{\ell})\geq v_{i}(g^{\prime}_{\ell+1})$. We have
@@
$\geq\sum_{\ell\in[k^{\prime}-1]}v_{i}(g^{\prime}_{\ell+1})+v_{i}(i_{n^{\prime}})=v_{i}(B_{i_{n^{\prime}}})\geq 1.\qed$
@@
returns a simple $\alpha$-MMS allocation with at most $\lfloor n^{*}/2\rfloor$ many distinct copies.
@@
Let $n$ be the number of agents after applying $\mathtt{S-reduce(\mathcal{I},\alpha)}$. Since $S^{\alpha}$ and $R^{\alpha}_{k}$ are valid reductions for all $k<|M|/n^{*}$, all agents who receive a bag during the $\mathtt{S-reduce(\mathcal{I},\alpha)}$ receive at least $\alpha$ fraction of their MMS value and the MMS value of the remaining agents do not decrease after this process. Without loss of generality, let the MMS value of the remaining agents be $1$ at this point of the algorithm. All the agents who receive a bag during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}$ get at least $\alpha$. Let agent $i$ be an agent who receives bag $B_{j}$ after the round robin phase. Let $g$ be the good that was duplicated and added to $B_{j}$ right before allocating it to agent $i$. $g\leq n^{\prime}$ and thus $v_{i}(g)\geq v_{i}(n^{\prime})$. By Lemma 18, $v_{i}(B_{j}\cup\{g\})\geq\alpha$. Therefore, the final allocation is $\alpha$-MMS.
We now prove that the generated MMS allocation is simple, namely that there are $t\leq\lfloor n^{*}/2\rfloor$ distinct copies, and that the $2t$ instances, composed of these copies and their corresponding original goods, are allocated to $2t$ distinct agents. Let $k=n^{*}-n$ be the number of agents who receive a bag in $\mathtt{S-reduce(\mathcal{I},\alpha)}$. Only when applying $S^{\alpha}$, new duplications are introduced and one duplication per two agents that are removed. So during $\mathtt{S-reduce(\mathcal{I},\alpha)}$, at most $\lfloor k/2\rfloor$ distinct copies are used. Each instance of these good is allocated to a different agent. Furthermore, all the instances of the goods that are allocated during $S^{\alpha}$ along with their owners, are removed from the instance. Therefore, these goods cannot be copied more than once and their owners cannot get any other goods later on. In the duplication phase of $\mathsf{BagFill\mbox{-}RoundRobin}$, we duplicate items $1,2,\ldots,n^{\prime}$ once and allocate the original and the duplications to $2n^{\prime}$ different agents. By Corollary 6, $n^{\prime}\leq\lfloor n/2\rfloor$. Therefore, in total at most $\lfloor n^{*}/2\rfloor$ distinct copies are used and the allocation is simple. ∎